Exhibit 12-1

(1)

(2)

(3)

(4)

Units of Factor X

Quantity of Output

Product Price

Marginal Revenue Product

0

10

$5

1

19

$5

A

2

27

$5

B

3

34

$5

C

4

39

$5

D


Refer to Exhibit 12-1. What value goes in blank C?

$35.00

$170.00

$40.00

$11.67
Answer #1

Marginal physical product = Q (3) - Q (2) = 34 - 27 = 7

MRP = Marginal physical product x Price = 7 x 5 = $ 35